OpenStudy (anonymous):

simplify the rational expression: x^3-1/1-x

OpenStudy (mertsj):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (mertsj):

Use that to factor the numerator and then use this:

OpenStudy (mertsj):

\[\frac{a-b}{b-a}=-1\]